Circular No. 7184

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V382 VELORUM S. Nakano, Sumoto, Japan, reports that T. Furuta, Kagiya-cho, Tokai,

recorded a prediscovery image of V382 Vel = N Vel 1999 on Fuji Color Super G400 film taken during his travel in New Zealand. His films yield the following magnitude estimates: May 20.57 UT, [13 (30-min exposure with 50-mm f/2.8 camera lens); 21.64, 4.3 (30-min exposure with 16-mm f/2.8 fish-eye lens). B. Sumner reports that M. Rudd, Mount Martha, Victoria, has also identified a prediscovery image of this nova (at mag 3.6 from comparison with ϕ Vel and π Cen, according to Sumner) on a 5-min exposure taken with Kodak Gold Flexiclear 400 film (28-mm camera lens) on May 21.715. Visual magnitude estimates, some communicated via T. A. Napoleao, São Paulo, Brazil: May 25.939 UT, 3.8 (J. G. de S. Aguiar, Campinas, Brazil); 26.25, 3.9 (S. J. O’Meara, Volcano, HI); 27.542, 4.1 (A. Pearce, Nedlands, W. Australia); 28.024, 4.5 (A. Alves, Florianopolis, Brazil); 28.41, 4.7 (R. H. McNaught, Bugaldie, N.S.W.); 29.990, 4.8 (Napoleao); 30.442, 4.7 (Pearce); 31.344, 5.2 (D. A. J. Seargent, The Entrance, N.S.W.); June 1.163, 5.2 (J. Garcia, Rama Caida, Argentina).

XTE J1550–564 R. Soria, Research School of Astronomy and Astrophysics, Australian

National University; and K. Wu, Special Research Centre for Theoretical Astrophysics, University of Sydney, write: “We have obtained a highresolution spectrum (FWHM = 0.12 nm) of the optical counterpart of the x-ray transient XTE J1550–564 on Apr. 10 with the Anglo-Australian Telescope at Siding Spring Observatory (SSO), and B, V, R, and I images on Apr. 22 and May 23 with the SSO 1.02-m telescope. The optical spectrum is now dominated by stellar features typical of a G6-G8 IV star: we detect narrow absorption lines at Hα (EW = 0.1 nm, FWHM = 50 km/s) and Hβ, while no emission is detected from He II 468.6-nm and the Bowen blend. We measure EW = 0.023 ± 0.002 nm and 0.18 ± 0.01 nm for the interstellar lines at 661.3 and 628.4 nm, respectively, and we estimate E(B − V) = 1.0 ± 0.1 (cf. IAUC 7113). Our photometric measurements on May 23.5 UT yield B = 21.83 ± 0.15, V = 20.02 ± 0.04, R = 18.97 ± 0.04, I = 17.68 ± 0.04. If the Roche-lobe-filling secondary star of XTE J1550–564 is a G8 subgiant, it has a mass of 1.20 ± 0.05 M_{⊙}, mean density of 0.06 ρ_{⊙}, and absolute magnitude M_{V} = 3.15 ± 0.05. Consequently we suggest a distance of 6.0 ± 0.7 kpc for the system, and a binary period of 1.5 ± 0.1 days. More optical observations of this blackhole candidate are encouraged, now that the x-ray source is in the ‘off’ state.”

1999 June 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green